# RotBench: Evaluating Multimodal Large Language Models on Identifying Image Rotation

### v2

@@ Abstract
We investigate to what extent Multimodal Large Language Models (MLLMs) can accurately identify the orientation of input images rotated 0°, 90°, 180°, and 270°. This task demands robust visual reasoning capabilities to detect rotational cues and contextualize spatial relationships within images, regardless of their orientation. To evaluate MLLMs on these abilities, we introduce RotBench – a 350-image manually-filtered benchmark comprising lifestyle, portrait, and landscape images. Despite the relatively simple nature of this task, we show that several state-of-the-art open and proprietary MLLMs, including GPT-5, o3, and Gemini-2.5-Pro, do not reliably identify rotation in input images. Providing models with auxiliary information – including captions, depth maps, and more – or using chain-of-thought prompting offers only small and inconsistent improvements. Our results indicate that most models are able to reliably identify right-side-up (0°) images, while certain models are able to identify upside-down (180°) images. None can reliably distinguish between 90° and 270°. Simultaneously showing the image rotated in different orientations leads to moderate performance gains for reasoning models, while a modified setup using voting improves the performance of weaker models. We further show that fine-tuning does not improve models’ ability to distinguish 90° and 270° rotations, despite substantially improving the identification of 180° images. Together, these results reveal a significant gap between MLLMs’ spatial reasoning capabilities and human perception in identifying rotation. [^1] Code and data: https://github.com/tianyiniu/RotBench
@@
Tianyi Niu  Jaemin Cho  Elias Stengel-Eskin  Mohit Bansal UNC Chapel Hill {tianyin4, jmincho, esteng, mbansal}@cs.unc.edu
@@ 1 Introduction
Figure 2: RotBench evaluation pipeline: for each image in RotBench, we rotate the image 0°, 90°, 180°, and 270° counter-clockwise. We represent the rotation estimation problem as a multiple-choice question answering problem (Section˜L.5), and separately measure accuracy on each image orientation. We optionally provide different forms of auxiliary information to aid the model in identifying image rotation. We emphasize that all forms of auxiliary information are separately extracted for each rotation; the ground truth rotation is not marked.
@@
Advancements in Multimodal Large Language Models (MLLMs) have led to increased performance in complex visual tasks, such as image-text retrieval, image segmentation, and visual question answering Li et al. (2025); Chen et al. (2023a); Ravi et al. (2024); Fu et al. (2024a); Chen et al. (2023b); Bai et al. (2025); OpenAI (2025c); Gemini Team (2025); Liu et al. (2023a); Deitke et al. (2024). However, a growing body of recent work suggests that MLLMs are sensitive to simple image transformations Anis et al. (2025), such as rotations, flips, and blurs, and they fail on tasks that are intuitive to humans Fu et al. (2024b); Pothiraj et al. (2025); Tong et al. (2024). Downstream tasks involving a rotating camera – such as robotic arm manipulation or first-person extreme sports analysis – require MLLMs to demonstrate robust spatial reasoning, regardless of image orientation. Given these challenges, this work explores a fundamental question: can MLLMs identify image orientation?
@@
Humans can quickly recognize whether an image has been rotated Shepard and Metzler (1971); Vandenberg and Kuse (1978); for example, it is easy for us to recognize that the left image in Fig.˜1 is not upright. The human viewer can use the orientation of the couch in the background to infer that the father is actually lying on his back rather than standing up. The simple task of identifying image rotation requires reconciling the image’s subjects, background, and semantics. Here, we show that identifying rotation remains a challenge, even in frontier MLLMs.
@@
We introduce RotBench (Section˜3), a benchmark for evaluating MLLMs’ ability to recognize rotation in images. RotBench consists of images sampled from Spatial-MM Shiri et al. (2024), and is comprised of two subsets: the 300-image RotBench-large and the 50-image RotBench-small. RotBench is carefully constructed to be challenging but fair, with a two-stage filtering procedure to remove images that are indistinguishable under different degrees of rotation. Section˜3 describes our dataset construction.
@@
Using RotBench, we explore whether frontier MLLMs can identify rotation in input images rotated 0°, 90°, 180°, and 270° (Fig.˜2). We also evaluate whether providing various forms of auxiliary information or using chain-of-thought Wei et al. (2022) prompting improves performance (Section˜4.3). We find that models are able to consistently identify right-side-up (0°) images. However, only stronger models are able to identify upside-down (180°) images. All models fail to accurately distinguish between 90° and 270° images (Section˜6.1, Section˜6.2). We find adding auxiliary information offers minimal and inconsistent improvement, often improving performance on 270° images at the expense of 90° images.
@@
Leveraging MLLMs’ tendency to accurately identify 0° images, we attempt to improve performance by further rotating an input image 0°, 90°, 180°, 270° and presenting MLLMs with all rotations simultaneously. Using this approach, reasoning models (o3 and Gemini-2.5-Pro) show performance improvements, while weaker models see performance degradation. Further extending this idea, we utilize a voting approach to algebraically obtain a majority vote for the correct ground truth orientation. We obtain a model prediction for each further rotation by subtracting the added angle. While this approach does show significant performance improvements on weaker models, it lacks scalability as it requires multiple model calls for each prediction, and assumes a priori knowledge of all possible orientations (Appendix˜I).
@@
Finally, fine-tuning on out-of-domain data (Section˜6.3) significantly improved performance on identifying 180° images but not 90° and 270°. Interestingly, we find an oscillating pattern of performance changes between 90° and 270° as training progresses. Any performance improvement in 90° is matched with a degradation of 270° and vice versa. These results suggest the presence of two local optima hindering across-the-board progress.
@@ 2 Related Work
Sensitivity to visual perturbations.  Previous work has shown that visual encoders and MLLMs are sensitive to simple image transformations. Anis et al. (2025) evaluate CLIP Radford et al. (2021) and SigLIP Zhai et al. (2023) on a suite of common image transformations – rotations, flips, noise, etc. – revealing substantial gaps between human and model understanding. Usama et al. (2025) finds MLLMs exhibit distinct failure patterns in scene-text and object reasoning tasks when applying ImageNet-C corruptions Hendrycks and Dietterich (2019) to image inputs. Taken together, these studies highlight that despite strong clean-image performance, visual encoders and MLLMs are highly sensitive to photometric and geometric distortions.
Image and camera orientation estimation.  Fine-tuning models to identify image orientation has been the focus of prior work Xu et al. (2024). For example, Fischer et al. (2015) and Joshi and Guerzhoy (2017) focused on fine-tuning convolutional neural networks (CNNs) to estimate and identify image rotation. While our work tackles a similar task, we are instead interested in the problem as a test of general-purpose MLLMs’ inherent reasoning abilities, i.e., whether they can estimate image rotation without extensive fine-tuning. Note that our work aligns more closely with the problem of image, not camera, orientation estimation. Camera orientation estimation is also a well-studied task in computer vision Xian et al. (2019). Instead of predicting the rotation of the image, camera orientation estimation seeks to predict the spatial location of the camera when capturing an image. Contemporary approaches of this task use deep networks to directly predict orientation parameters from image features in an end-to-end manner Xian et al. (2019); Lee et al. (2021, 2020).
@@
Gap between human perception and MLLMs.  A growing body of work shows MLLMs exhibit fundamental gaps compared to human perceptual capabilities. Pothiraj et al. (2025) propose CAPTURe – a benchmark for occluded object counting – and report sharp drops in model accuracy on both synthetic and real images. Zhou et al. (2025) proposes MMVM, a benchmark for visual matching across images. Fu et al. (2024b) collect BLINK, a dataset comprised of visual tasks humans can solve in a ‘blink,’ such as identifying visual similarity and relative depth. Both Zhou et al. (2025) and Fu et al. (2024b) report low zero-shot accuracy on their respective tasks, suggesting MLLMs lack many of the intuitive reasoning mechanisms that underpin human visual perception. In this vein, our work provides a novel perspective to analyze and interpret the spatial reasoning capabilities of MLLMs, with results indicating that models struggle with this previously underexplored challenge.
@@ 3 RotBench
We introduce RotBench, a benchmark for evaluating models’ ability to identify rotation in input images. RotBench is created using images from Spatial-MM Shiri et al. (2024) and includes two subsets: the 300-image RotBench-large and the 50-image RotBench-small. While rotating an image is straightforward, not all images are meaningful under rotation. A rotated portrait of a human, such as the image shown in Fig.˜2, will be easily noticed by human viewers. However, a top-down view of a simple tabletop does not significantly differ when rotated (Fig.˜9). We use a two-stage filtering process to ensure different rotations of each image are clearly distinguishable. This section provides an overview of our filtering procedure. Appendix˜A describes our dataset procedure and statistics in further detail.
Stage 1.  We randomly sample 300 images from Spatial-MM Shiri et al. (2024). Stage 1 involves a single annotator. Depending on the amount of visual signals available, the Stage 1 annotator decides to either accept, discard, or flag each image. Flagged images then proceed to Stage 2. We provide further examples and details of accepted, flagged, and discarded images in Fig.˜9.
Stage 2.  Stage 2 involves a group of three human evaluators. Each flagged image is rotated 0°, 90°, 180°, and 270° counter-clockwise then presented to the evaluators as multiple-choice questions (Section˜A.4). During Stage 2, an evaluator will see each image four times, once per orientation. Any image that elicits an incorrect answer from two or more evaluators across all four orientations is discarded. Otherwise, the image is accepted. Stage 1 and 2 are repeated until a total of 300 images are accepted.
RotBench-large and RotBench-small.  All images that have been accepted in Stage 1 and 2 are organized into RotBench-large. As each image is rotated in four orientations, obtaining human performance on RotBench-large is costly. Luckily, Stage 2 provides a human baseline for the subset of RotBench-large images that have been flagged (25 images). We expand this subset by further sampling images from Spatial-MM. From these additional images, we only select images that fit the criteria for flagging to proceed to another round of Stage 2 evaluation. This process repeats until we reach a total of 50 images, organizing them into RotBench-small. [^2] The human evaluators all exhibit high accuracy, averaging ¿ 0.97 for all rotations.
@@ 4.1 Models
We evaluate various open-weight and proprietary MLLMs on RotBench: Qwen-2.5-VL-7B-Instruct Bai et al. (2025); GPT-4o OpenAI et al. (2024), GPT-4.1 OpenAI (2025a), o3 OpenAI (2025c), GPT-5 OpenAI (2025b), Gemini-2.0-Flash Hassabis and Kavukcuoglu (2024), Gemini-2.5-Flash Gemini Team (2025), and Gemini-2.5-Pro Gemini Team (2025). Due to cost and resource limitations, we evaluate Gemini-2.5-Flash, Gemini-2.5-Pro, GPT-4.1, GPT-5, and o3 on RotBench-small. Responses are obtained through greedy decoding, while all chain-of-thought Wei et al. (2022) responses are obtained with a temperature of 0.3.333We perform an ablation study where we vary the sampling temperature (Table 8). We use the default sampling temperature for proprietary models that do not expose a temperature parameter in their API interface.
@@ 4.2 Setup and Evaluation
We rotate each image in RotBench-large and RotBench-small by 0°, 90°, 180°, and 270° counter-clockwise, resulting in a total of 1000 and 200 images. Note that a 90° counter-clockwise rotation is equivalent to a 270° clockwise rotation. For each image and orientation, we provide the model with the image, a brief description of the task, and various forms of auxiliary information (Section˜4.3). We frame this task as a four-way classification problem. To ensure robustness, the mapping between letter choice and degree of rotation is randomized for each prompt. We evaluate models on RotBench-large and report average accuracy and standard deviation across 3 runs in Table˜1. We use the same procedure, albeit with only 2 runs, on RotBench-small and report results in Table˜2. All prompts used are available in Appendix˜L.
@@ 4.3 Auxiliary Information
Figure˜3 illustrates all forms of auxiliary information provided to the model. Note that all auxiliary information is separately extracted for each rotation, ensuring our approach does not depend on prior knowledge of the image’s orientation.
Captions.  For each image and rotation, we instruct GPT-4o to provide a detailed caption (Section˜L.2). We emphasize that each image is captioned four times, once per rotation.
Bounding Boxes.  For each image and rotation, we first use GPT-4o to extract the primary subjects within the image (Section˜L.1). Next, along with the image, the list of subjects is given to GroundingDINO Liu et al. (2023b) to extract a set of normalized coordinates for each subject, [^4] Each set of coordinates is a four-element tuple, composed of [x_min,y_min,x_max,y_max][x\_min,y\_min,x\_max,y\_max]. which is directly injected into the prompt.
@@ 5 Main Results
Table˜1 displays the results of evaluating Qwen-2.5-VL-7B-Instruct, Gemini-2.0-Flash, and GPT-4o along with various auxiliary information on RotBench-large. Table˜2 displays results of evaluating Qwen-2.5-VL-7B-Instruct, GPT-4o, GPT-4.1, GPT-5, o3, Gemini-2.0-Flash, Gemini-2.5-Flash, Gemini-2.5-Pro on RotBench-small. Due to the high token cost of proprietary reasoning models, we only evaluate zero-shot and CoT prompts. However, we evaluate providing rotation grids to o3 and Gemini-2.5-Pro. [^5] Table 5 further shows results obtained from evaluating these two models on RotBench-small with all available auxiliary information.
@@
None of the auxiliary information results in meaningful, consistent performance gains across all tested models. Paradoxically, the introduction of additional information sometimes leads to marginal performance degradation. When including all forms of auxiliary information, Qwen’s accuracy on 90° decreased from 0.51 to 0.26. Similarly, Gemini-2.0-Flash’s accuracy on 270° decreased from 0.44 to 0.17.
@@
Providing the rotation grid degraded performance on most models. Gemini-2.0-Flash accuracy on 90° decreased by nearly 0.5 compared to CoT prompting. Rotation grid guided, however, improved GPT-4o accuracy on 270° by around 0.1. At the same time, both o3 and Gemini-2.5-Pro saw performance improvements. Notably, Gemini-2.5-Pro’s accuracy on 90° and 270° improved by 0.15. These results suggest the two reasoning models are much more effective at utilizing visual context. We further leverage the robust identification of 0° images and modify the rotation grid into a majority voting approach Appendix˜I. This setup results in gains even in weaker models, indicating MLLMs can achieve moderate performance with sufficient visual scaffolding.
@@
Employing chain-of-thought (CoT) prompting yields mixed results, differing notably across models and orientations. Gemini models show improved accuracy on 90° rotations but decreased accuracy for 270°. Conversely, GPT models exhibit improved performance on 270° rotations, coupled with degraded accuracy for 90° rotations. However, CoT consistently enhanced accuracy for 180° rotations across all models tested. These findings suggest CoT prompting can help models better reason about rotations to some extent, but does not universally resolve inherent challenges in distinguishing portrait orientations. Appendix˜F examines a common failure of CoT in detail.

### v3

@@ Abstract
We investigate to what extent Multimodal Large Language Models (MLLMs) can accurately identify the orientation of input images rotated 0°, 90°, 180°, and 270°. This task demands robust visual reasoning capabilities to detect rotational cues and contextualize spatial relationships within images, regardless of their orientation. To evaluate MLLMs on these abilities, we introduce RotBench, a 350-image manually-filtered benchmark comprising lifestyle, portrait, and landscape images. Despite the relatively simple nature of this task, we show that several state-of-the-art open and proprietary MLLMs, including GPT-5, o3, and Gemini-2.5-Pro, do not reliably identify rotation in input images. Providing models with auxiliary information—including captions, depth maps, and more—or using chain-of-thought prompting offers only small and inconsistent improvements. Our results indicate that most models are able to reliably identify right-side-up (0°) images, while certain models are able to identify upside-down (180°) images. None can reliably distinguish between 90° and 270° rotated images. Simultaneously showing the image rotated in different orientations leads to moderate performance gains for reasoning models, while a modified setup using voting improves the performance of weaker models. We further show that fine-tuning does not improve models’ ability to distinguish 90° and 270° rotations, despite substantially improving the identification of 180° images. Together, these results reveal a significant gap between MLLMs’ spatial reasoning capabilities and human perception in identifying rotation. [^1] Code and data: https://github.com/tianyiniu/RotBench
@@
Tianyi Niu1   Jaemin Cho2   Elias Stengel-Eskin3   Mohit Bansal1 1UNC Chapel Hill 2Allen Institute for Artificial Intelligence 3The University of Texas at Austin
@@ 1 Introduction
Figure 2: RotBench evaluation pipeline: for each image in RotBench, we rotate the image 0°, 90°, 180°, and 270° counter-clockwise. We represent the rotation estimation problem as a multiple-choice question answering problem (Appendix˜P), and separately measure accuracy on each image orientation. We optionally provide different forms of auxiliary information to aid the model in identifying image rotation. We emphasize that all forms of auxiliary information are separately extracted for each rotation; the ground truth rotation is not marked.
@@
Advancements in Multimodal Large Language Models (MLLMs) have led to increased performance in complex visual tasks, such as image-text retrieval, image segmentation, and visual question answering Li et al. (2025); Chen et al. (2023a); Ravi et al. (2024); Fu et al. (2024a); Chen et al. (2023b); Bai et al. (2025); OpenAI (2025c); Gemini Team (2025); Liu et al. (2023a); Deitke et al. (2024a). However, a growing body of recent work suggests that MLLMs are sensitive to simple image transformations Anis et al. (2025), such as rotations, flips, and blurs, and they fail on tasks that are intuitive to humans Fu et al. (2024b); Pothiraj et al. (2025); Tong et al. (2024). Downstream tasks involving a rotating camera, such as robotic arm manipulation or first-person extreme sports analysis, require MLLMs to demonstrate robust spatial reasoning, regardless of image orientation (Appendix˜C). Given these challenges, this work explores a fundamental question: can MLLMs identify image orientation?
@@
Humans can quickly recognize whether an image has been rotated Shepard and Metzler (1971); Vandenberg and Kuse (1978); for example, it is easy for us to recognize that the left image in Fig.˜1 is not upright. A human viewer can use the orientation of the couch in the background to infer that the father is actually lying on his back rather than standing up. The simple task of identifying image rotation requires reconciling the image’s subjects, background, and semantics. Here, we show that identifying rotation remains a challenge, even in frontier MLLMs.
@@
We introduce RotBench (Section˜3), a benchmark for evaluating MLLMs’ ability to recognize rotation in images. RotBench consists of images sampled from Spatial-MM Shiri et al. (2024) and is comprised of two subsets: the 300-image RotBench-large and the 50-image RotBench-small. RotBench is carefully constructed to be challenging but fair, with a two-stage filtering procedure to remove images that are indistinguishable under different degrees of rotation. Section˜3 describes our dataset construction.
@@
Using RotBench, we explore whether frontier MLLMs can identify rotation in input images rotated 0°, 90°, 180°, and 270° (Fig.˜2). We also evaluate whether providing various forms of auxiliary information or using chain-of-thought Wei et al. (2022) prompting improves performance (Section˜4.3). We find that models are able to consistently identify right-side-up (0°) images. However, only stronger models are able to identify upside-down (180°) images. All models fail to accurately distinguish between 90° and 270° images (Section˜5, Section˜6.1, Section˜6.2). We find adding auxiliary information offers minimal and inconsistent improvement, often improving performance on 270° images at the expense of 90° images.
@@
Leveraging MLLMs’ tendency to accurately identify 0° images, we attempt to improve performance by further rotating an input image 0°, 90°, 180°, 270° and presenting MLLMs with all rotations simultaneously (Section˜5). Using this approach, reasoning models show performance improvements, while weaker models see performance degradation. Further extending this idea, we utilize a voting approach to algebraically obtain a majority vote for the correct ground truth orientation. We obtain a model prediction for each further rotation by subtracting the added angle. While this approach does show significant performance improvements on weaker models, it lacks scalability as it requires multiple model calls for each prediction, and assumes a priori knowledge of all possible orientations (Section˜6.4).
@@
Finally, fine-tuning on out-of-domain data (Section˜6.3) significantly improved performance on identifying 180° images but not 90° and 270°. Interestingly, we find an oscillating pattern of performance changes between 90° and 270° as training progresses. Any performance improvement in 90° is matched with a degradation in performance of 270° and vice versa. These results suggest the presence of two local optima hindering across-the-board progress.
@@ 2 Related Work
This section reviews the literature most closely related to our work. Additional relevant areas are discussed in Appendix˜B.
Image orientation estimation.  Fine-tuning models to identify image orientation has been the focus of prior work Xu et al. (2024). For example, Fischer et al. (2015) and Joshi and Guerzhoy (2017) focused on fine-tuning convolutional neural networks (CNNs) to estimate and identify image rotation. While our work tackles a similar task, we are instead interested in the problem as a test of general-purpose MLLMs’ inherent reasoning abilities, i.e., whether they can estimate image rotation without extensive fine-tuning. Note that our work aligns more closely with the problem of image, not camera, orientation estimation. We further discuss camera orientation estimation in Appendix˜B.
@@
Gap between human perception and MLLMs.  A growing body of work shows MLLMs exhibit fundamental gaps compared to human perceptual capabilities. Pothiraj et al. (2025) propose CAPTURe, a benchmark for occluded object counting, and report sharp drops in model accuracy on both synthetic and real images. Zhou et al. (2025) proposes MMVM, a benchmark for visual matching across images. Fu et al. (2024b) collect BLINK, a dataset comprised of visual tasks humans can solve in a ‘blink,’ such as identifying visual similarity and relative depth. Both Zhou et al. (2025) and Fu et al. (2024b) report low zero-shot accuracy on their respective tasks, suggesting MLLMs lack many of the intuitive reasoning mechanisms that underpin human visual perception. In this vein, our work provides a novel perspective to analyze and interpret the spatial reasoning capabilities of MLLMs, with results indicating that models struggle with this previously underexplored challenge.
@@ 3 RotBench
We introduce RotBench, a benchmark for evaluating models’ ability to identify rotation in input images. RotBench is created using images from Spatial-MM Shiri et al. (2024) and includes two subsets: the 300-image RotBench-large and the 50-image RotBench-small. While rotating an image is straightforward, not all images are meaningful under rotation. A rotated portrait of a human, such as the image shown in Fig.˜2, will be easily noticed by human viewers. However, a top-down view of a simple tabletop does not significantly differ when rotated (Fig.˜7). We use a two-stage filtering process to ensure different rotations of each image are clearly distinguishable. This section provides an overview of our filtering procedure. Appendix˜A describes our dataset procedure and statistics in further detail.
Stage 1.  We randomly sample 300 images from Spatial-MM Shiri et al. (2024). Stage 1 involves a single annotator. Depending on the amount of visual signals available, the Stage 1 annotator decides to either accept, discard, or flag each image. Flagged images then proceed to Stage 2. We provide further examples and details of accepted, flagged, and discarded images in Fig.˜7.
Stage 2.  Stage 2 involves a group of three human evaluators. Each flagged image is rotated 0°, 90°, 180°, and 270° counter-clockwise and then presented to the evaluators as multiple-choice questions (Section˜A.4). During Stage 2, an evaluator will see each image four times, once per orientation. Any image that elicits an incorrect answer from two or more evaluators across all four orientations is discarded. Otherwise, the image is accepted. Stages 1 and 2 are repeated until 300 images are accepted.
RotBench-large and RotBench-small.  All images that have been accepted in Stages 1 and 2 are organized into RotBench-large. As each image is rotated in four orientations, obtaining human performance on RotBench-large is costly. Fortunately, Stage 2 provides a human baseline for the subset of RotBench-large images that have been flagged (25 images). We expand this subset by further sampling images from Spatial-MM. From these additional images, we only select images that fit the criteria for flagging to proceed to another round of Stage 2 evaluation. This process repeats until we reach a total of 50 images, organizing them into RotBench-small. [^2] The human evaluators all exhibit high accuracy, averaging ¿ 0.97 for all rotations.
@@ 4.1 Models
We evaluate various open-weight and proprietary MLLMs on RotBench: Qwen-2.5-VL-7B-Instruct Bai et al. (2025); GPT-4o OpenAI et al. (2024), GPT-4.1 OpenAI (2025a), o3 OpenAI (2025c), GPT-5 OpenAI (2025b), Gemini-2.0-Flash Hassabis and Kavukcuoglu (2024), Gemini-2.5-Flash Gemini Team (2025), and Gemini-2.5-Pro Gemini Team (2025). Due to cost and resource limitations, we evaluate Gemini-2.5-Flash, Gemini-2.5-Pro, GPT-4.1, GPT-5, and o3 on RotBench-small. Responses are obtained through greedy decoding, while all chain-of-thought Wei et al. (2022) responses are obtained with a temperature of 0.3.333We perform an ablation study where we vary the sampling temperature (Table 10). We use the default sampling temperature for proprietary models that do not expose a temperature parameter in their API interface. In addition, we provide further results on Llama-3.2-11B-Instruct Grattafiori et al. (2024), Qwen-2.5-VL-3B-Instruct Bai et al. (2025), Phi-4-MM-Instruct Microsoft et al. (2025), Gemma-3-12B-Instruct Gemma et al. (2025), Molmo-7B-O Deitke et al. (2024b), and Claude-4.1-Opus Anthropic (2025) in Appendix˜E.
@@ 4.2 Setup and Evaluation
We rotate each image in RotBench-large and RotBench-small by 0°, 90°, 180°, and 270° counter-clockwise, resulting in a total of 1,200 and 200 images. Note that a 90° counter-clockwise rotation is equivalent to a 270° clockwise rotation. For each image and orientation, we provide the model with the image, a brief description of the task, and various forms of auxiliary information (Section˜4.3). We frame this task as a four-way classification problem. To ensure robustness, the mapping between letter choice and degree of rotation is randomized for each prompt. We evaluate models on RotBench-large and report average accuracy and standard deviation across 3 runs in Table˜1. We use the same procedure, albeit with only 2 runs, on RotBench-small and report results in Table˜2. All prompts used are given in Appendix˜P. We also present evaluations under a regression formulation and a multi-choice setup with increased granularity in Appendix˜K, all leading to inferior performance. We select our current setup for its simplicity and because it already challenges frontier models.
@@ 4.3 Auxiliary Information
Figure˜9 illustrates all forms of auxiliary information provided to the model. Note that all auxiliary information is separately extracted for each rotation, ensuring our approach does not depend on prior knowledge of the image’s orientation.
Captions.  For each image and rotation, we instruct GPT-4o to provide a detailed caption (Appendix˜P). We emphasize that each image is captioned four times, once per rotation.
Bounding Boxes.  For each image and rotation, we first use GPT-4o to extract the primary subjects within the image (Appendix˜P). Next, along with the image, the list of subjects is given to GroundingDINO Liu et al. (2023b) to extract a set of normalized coordinates for each subject, [^4] Each set of coordinates is a four-element tuple, composed of [x_min, y_min, x_max, y_max]. which is directly injected into the prompt.
@@ 5 Main Results
Table˜1 displays the results of evaluating Qwen-2.5-VL-7B-Instruct, Gemini-2.0-Flash, and GPT-4o along with various auxiliary information on RotBench-large. Table˜2 displays results of evaluating Qwen-2.5-VL-7B-Instruct, GPT-4o, GPT-4.1, GPT-5, o3, Gemini-2.0-Flash, Gemini-2.5-Flash, Gemini-2.5-Pro on RotBench-small. Due to the high token cost of proprietary reasoning models, we only evaluate zero-shot and CoT prompts. However, we evaluate providing rotation grids to o3 and Gemini-2.5-Pro. [^5] Table 6 further shows results obtained from evaluating these two models on RotBench-small with all available auxiliary information. We provide qualitative examples of images that models answer correct and incorrectly in Appendix˜N.
@@
None of the auxiliary information results in meaningful, consistent performance gains across all tested models. Paradoxically, the introduction of additional information sometimes leads to marginal performance degradation. When including all forms of auxiliary information, Qwen’s accuracy on 90° decreased from 0.51 to 0.26. Similarly, Gemini-2.0-Flash’s accuracy on 270° decreased from 0.44 to 0.17. We further examine why scene graphs fails to consistently improve model performance in Appendix˜J.
@@
Providing the rotation grid degraded performance on most models. Gemini-2.0-Flash accuracy on 90° decreased by nearly 0.5 compared to CoT prompting. Rotation grid guided, however, improved GPT-4o accuracy on 270° by around 0.1. At the same time, both o3 and Gemini-2.5-Pro saw performance improvements. Notably, Gemini-2.5-Pro’s accuracy on 90° and 270° improved by 0.15. These results suggest the two reasoning models are much more effective at utilizing visual context. We further leverage the robust identification of 0° images and modify the rotation grid into a majority voting approach (Section˜6.4). This setup results in gains even in weaker models, indicating MLLMs can achieve moderate performance with sufficient visual scaffolding.
@@
Employing chain-of-thought (CoT) prompting yields mixed results. Gemini models show improved accuracy on 90° rotations but decreased accuracy for 270°. On the other hand, GPT models displayed inverse trends. Yet, CoT consistently enhanced accuracy for 180° rotations across all models tested. These findings suggest CoT prompting can help models better reason about rotations to some extent, but does not universally resolve inherent challenges in distinguishing portrait orientations. Appendix˜H closely examines a common failure of CoT.